# Optimization of Prompt Learning via Multi-Knowledge Representation for Vision-Language Models

### v2

@@ 4.3 Ablation Study
Position of Classname Varies.  In the context of optimization, the position of the classname can be either at the end or in the middle [36]. Setting $\beta=0.6$, optimizing the context length to 16 and PK as the input. We explored the impact of different classname positions on experimental results across 11 datasets. The outcomes are presented in Table 6. Overall performance is better when the classname position is at the end rather than in the middle.
@@
Different Context Lengths.  We set the context lengths to 8, 16, and 32, with PK as the input and the $\beta$ value set to 0.6. We evaluated the mean comparison after training with 16 shots on 11 datasets to explore the impact of different context lengths on the results. The results are shown in Table 7. Our method demonstrates robust performance across varying context lengths, with minimal perturbation to the outcomes attributable to changes in context length.
5 CONCLUSION
@@
In this paper, we have gained insights into the importance of Multi-Knowledge Representation for VLMs. Existing technologies have overlooked the method of context template learning, which incorporates domain knowledge. Such oversight will restrict VLM’s ability to learn Multi-Knowledge Representation during the pre-training phase, further limiting its performance in downstream tasks. To address this issue, we propose Context Optimization with Multi-Knowledge Representation (CoKnow). We trained lightweight semantic knowledge mappers, capable of generating Multi-Knowledge Representation without requiring additional inputs, while also constructing a knowledge-guided Prompt learning framework. We extensively evaluated it on multiple benchmark tests, and the results indicate that the proposed CoKnow is an effective method for prompt learning.
Appendix A Multi-Knowledge Generation
In this section, we offer detailed methodologies and concrete exemplars of Multi-Knowledge generation.
This paper investigates efficacious methodologies for Multi-Knowledge Generation across 14 datasets encompassing diverse fields. (Table 8 provides a summary of these 14 datasets.) The approach entails furnishing GPT-4 with scalable prompts to elicit its responses, thereby facilitating the acquisition of Multi-Knowledge. In Table LABEL:tab:long, we furnish the prompts to generate various types of Multi-Knowledge across the 14 datasets, alongside the Hand-crafted Templates employed for these datasets.
Table 8: Datasets statistics.
Dataset Classes Train Val Test
ImageNet 1,000 1.28M N/A 50,000
Caltech101 100 4,128 1,649 2,465
OxfordPets 37 2,944 736 3,669
StanfordCars 196 6,509 1,635 8,041
Flowers102 102 4,093 1,633 2,463
Food101 101 50,500 20,200 30,300
FGVCAircraft 100 3,334 3,333 3,333
SUN397 397 15,880 3,970 19,850
DTD 47 2,820 1,128 1,692
EuroSAT 10 13,500 5,400 8,100
UCF101 101 7,639 1,898 3,783
CIFAR-10 10 50,000 N/A 10,000
CIFAR-100 10 50,000 N/A 10,000
STL-10 10 5,000 N/A 8,000
In Figure 7, we present a specific instance to generate PK utilizing GPT-4. Upon providing a category face to GPT-4 as input, the comprehensive prompt yields the following response: Area with eyes, nose, mouth; center of emotional expression.
Figure 7: The specific process of generating Multi-Knowledge.
Appendix B Instances of Multi-Knowledge
In this section, we provide detailed instances of Visual Knowledge (VK), Non-Visual Knowledge (NVK), and Panoramic Knowledge (PK). As shown in Figure 8, we randomly select multiple images and provide Multi-Knowledge generated based on their class names.
Figure 8: Instances of Multi-Knowledge.
Appendix C Experiments of zero-shot prediction with Multi-knowledge
To explore the significance of Multi-Knowledge in pre-trained Vision-Language Models (VLMs), we conducted zero-shot experiments on the CIFAR-10, CIFAR-100, and STL10 datasets. We utilize Vision Transformer (ViT) as the visual encoding backbone for CLIP and perform experiments on the CIFAR-10 test set, comprising 10,000 images. Our prediction process is illustrated in Figure 9 (which is identical to the method described in Figure 1 of the main text). In Figure 9 (b), we use the PK, VK, and NVK generated by GPT-4 as inputs to the text encoder, to obtain the Multi-Knowledge Representation corresponding to the image.
Figure 9: The prediction process in CLIP before and after the introduction of Multi-Knowledge.
In Figure 10, we present the zero-shot prediction with the incorporation of various types of Multi-Knowledge. Notably, PK yields the most substantial enhancement in accuracy. These experimental findings underscore the significance of Multi-Knowledge. As illustrated in Figure 11, we conduct a t-SNE visualization of the combined vector results from the original image representation and the Multi-Knowledge Representation. It can be observed that on various datasets, the representations are clearly segmented into multiple clusters, which correspond to the number of categories.
Figure 10: The prediction results with the introduction of different Multi-Knowledge. Among these results, PK achieves the best performance.
Table 9: The prompts given to GPT-4 for Multi-Knowledge Generation.
Dataset VK NVK PK Hand-crafted Template
Imagenet Based on the category name, please provide the most distinctive visual features unique to this category. Requirements: 1. Do not use the same appearance description as other categories. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid mentioning colors. Based on this category name, please provide the most unique feature that transcends visual concepts and distinguishes it from other categories, without visual descriptions. Requirements: 1. Do not use the same description as other categories. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid visual descriptions. Based on this category’s name, provide the most distinguishing feature description. Requirements: 1. Do not use the same description as other categories. 2. Keep it concise, no more than 10 words. 3. The description must only include visual and higher concept features of this category. 4. Ensure clarity, no ambiguity. 5. Do not describe color. 6. Must include visual feature of the category. 7. Must include higher concept feature of the category. a photo of a [CLASS].
Eurosat This is the category name of images captured by satellites. Please describe the most distinctive visual feature unique to this category. Requirements: 1. Do not use the same appearance description as other categories. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid mentioning colors. This is the category name of images captured by satellites. Please provide the most unique feature that transcends visual concepts and distinguishes it from other categories, without visual descriptions. Requirements: 1. Do not use the same description as other categories. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid visual descriptions. This is the category name of images captured by satellites. Please provide the most distinguishing feature that transcends visual concepts from other categories. Requirements: 1. Do not use the same description as the satellite image category. 2. Keep it concise, no more than 10 words. 3. The description must only include visual and higher concept features of this category. 4. Ensure clarity, no ambiguity. 5. Do not describe color. 6. Must include the visual feature of this category. 7. Must include higher concept feature of this category. a centered satellite photo of [CLASS].
Caltech101 Based on the category name, please provide the most distinctive visual features unique to this category. Requirements: 1. Do not use the same appearance description as other categories. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid mentioning colors. Based on this category name, please provide the most unique feature that transcends visual concepts and distinguishes it from other categories, without visual descriptions. Requirements: 1. Do not use the same description as other categories. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid visual descriptions. Based on this category’s name, provide the most distinguishing feature description. Requirements: 1. Do not use the same description as other categories. 2. Keep it concise, no more than 10 words. 3. The description must only include visual and higher concept features of this category. 4. Ensure clarity, no ambiguity. 5. Do not describe color. 6. Must include visual feature of the category. 7. Must include higher concept feature of the category. a photo of a [CLASS].
FGVCaircraft This is the category name of a type of aircraft model. Please provide a description of the most distinctive visual features unique to this category. Requirements: 1. Do not use the same appearance description as other aircraft models. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid mentioning colors. This is the category name of a type of aircraft model. Please provide the most unique feature that transcends visual concepts and distinguishes it from other categories, without visual descriptions. Requirements: 1. Do not use the same description as other aircraft models. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid visual descriptions. This is a category of aircraft models. Please provide the most distinguishing feature from other aircraft model categories. Requirements: 1. Do not use the same description as other aircraft model categories. 2. Keep it concise, no more than 10 words. 3. The description must only include visual and higher concept features of this category. 4. Ensure clarity, no ambiguity. 5. Do not describe color. 6. Must include the visual feature of this category. 7. Must include higher concept feature of this category. a photo of a [CLASS],a type of aircraft.
Sun397 This is the name of a real-life scene category. Please provide a description of the most distinctive visual features unique to this category. Requirements: 1. Do not use the same appearance description as other categories. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid mentioning colors. This is the name of a real-life scene category. Please provide the most distinctive feature that transcends visual concepts and distinguishes it from other categories, without visual descriptions. Requirements: 1. Do not use the same description as other real-life scene categories. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid visual descriptions. This is the category name of a type of real-life scene. Please provide a description of the most distinctive feature that sets it apart from other scene categories. Requirements: 1. Do not use the same description as other scene categories. 2. Keep it concise, no more than 10 words. 3. The description must only include visual and higher conceptual features of this category. 4. Ensure clarity without ambiguity. 5. Avoid mentioning colors. 6. Must include visual features of this category. 7. Must include higher conceptual features of this category. a photo of a [CLASS].
Oxford pets This is the name of an animal species. Please provide a description of the most distinctive visual features unique to this category. Requirements: 1. Do not use the same appearance description as other species. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid mentioning colors. This is the name of a pet breed. Please provide the most distinctive feature that transcends visual concepts and distinguishes it from other breeds, without visual descriptions. Requirements: 1. Do not use the same description as other pet breeds. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid visual descriptions. This is the category name of a type of pet breed. Please provide a description of the most distinctive feature that sets it apart from other pet breeds. Requirements: 1. Do not use the same description as other pet breed categories. 2. Keep it concise, no more than 10 words. 3. The description must only include visual and higher conceptual features of this category. 4. Ensure clarity without ambiguity. 5. Avoid mentioning colors. 6. Must include visual features of this category. 7. Must include higher conceptual features of this category. a photo of a [CLASS], a type of pet.
Stanford cars This is the name of a category of cars. Please provide a description of the most distinctive visual features unique to this category. Requirements: 1. Do not use the same appearance description as other categories. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid mentioning colors. This is the name of a category of car brands. Please provide the most distinctive feature that transcends visual concepts and distinguishes it from other categories, without visual descriptions. Requirements: 1. Do not use the same description as other car brand categories. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid visual descriptions. This is the category name of a type of car. Please provide a description of the most distinctive feature that sets it apart from other car categories. Requirements: 1. Do not use the same description as other car categories. 2. Keep it concise, no more than 10 words. 3. The description must only include visual and higher conceptual features of this category. 4. Ensure clarity without ambiguity. 5. Avoid mentioning colors. 6. Must include visual features of this category. 7. Must include higher conceptual features of this category. a photo of a [CLASS].
Cifar10 Based on the category name, please provide the most distinctive visual features unique to this category. Requirements: 1. Do not use the same appearance description as other categories. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid mentioning colors. Based on this category name, please provide the most unique feature that transcends visual concepts and distinguishes it from other categories, without visual descriptions. Requirements: 1. Do not use the same description as other categories. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid visual descriptions. Based on this category’s name, provide the most distinguishing feature description. Requirements: 1. Do not use the same description as other categories. 2. Keep it concise, no more than 10 words. 3. The description must only include visual and higher concept features of this category. 4. Ensure clarity, no ambiguity. 5. Do not describe color. 6. Must include visual feature of the category. 7. Must include higher concept feature of the category. a photo of a [CLASS].
DTD This is the name of a texture category. Please provide a description of the most distinctive visual features unique to this category. Requirements: 1. Do not use the same appearance description as other categories. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid mentioning colors. This is the name of a texture category. Please provide the most unique feature that transcends visual concepts and distinguishes it from other categories, without visual descriptions. Requirements: 1. Do not use the same description as other texture categories. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid visual descriptions. This is the category name of a type of texture. Please provide a description of the most distinctive feature that sets it apart from other texture categories. Requirements: 1. Do not use the same description as other texture categories. 2. Keep it concise, no more than 10 words. 3. The description must only include visual and higher conceptual features of this category. 4. Ensure clarity without ambiguity. 5. Avoid mentioning colors. 6. Must include visual features of this category. 7. Must include higher conceptual features of this category. [CLASS] texture.
Food101 This is a category of food. Please provide the most distinctive visual feature of this category. Requirements: 1. Do not use the same appearance description as other categories. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid mentioning colors. This is the category name of a type of food. Please provide the most distinctive feature that transcends visual concepts and distinguishes it from other categories, without visual descriptions. Requirements: 1. Do not use the same description as other categories. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid visual descriptions. This is the name of a category of food. Please provide the most distinctive feature that sets it apart from other food categories. Requirements: 1. Do not use the same description as other food categories. 2. Keep it concise, no more than 10 words. 3. The description must only include visual and higher conceptual features of this category. 4. Ensure clarity without ambiguity. 5. Do not mention colors. 6. Must include visual features of this category. 7. Must include higher conceptual features of this category. a photo of [CLASS], a type of food.
UCF101 This is the name of a character action behavior. Please provide a description of the most distinctive visual features unique to this action. Requirements: 1. Do not use the same appearance description as other actions. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid mentioning colors. This is the name of a category of human behavior actions. Please provide the most distinctive feature that transcends visual concepts and distinguishes it from other categories, without visual descriptions. Requirements: 1. Do not use the same description as other categories. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid visual descriptions. This is the name of a human behavior action. Please provide a description of the most distinctive feature that sets it apart from other action behavior categories. Requirements: 1. Do not use the same description as other behavior categories. 2. Keep it concise, no more than 10 words. 3. The description must only include visual and higher conceptual features of this category. 4. Ensure clarity without ambiguity. 5. Avoid mentioning colors. 6. Must include visual features of this category. 7. Must include higher conceptual features of this category. a photo of a person doing [CLASS].
Oxford flowers This is the name of a category of flowers. Please provide a description of the most distinctive visual features unique to this category. Requirements: 1. Do not use the same appearance description as other flower categories. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid mentioning colors. This is the name of a category of flowers. Please provide the most distinctive feature that transcends visual concepts and distinguishes it from other categories, without visual descriptions. Requirements: 1. Do not use the same description as other flower categories. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid visual descriptions. This is the name of a category of flowers. Please provide a description of the most distinctive feature that sets it apart from other flower categories. Requirements: 1. Do not use the same description as other flower categories. 2. Keep it concise, no more than 10 words. 3. The description must only include visual and higher conceptual features of this category. 4. Ensure clarity without ambiguity. 5. Avoid mentioning colors. 6. Must include visual features of this category. 7. Must include higher conceptual features of this category. a photo of a [CLASS], a type of flower.
Cifar100 Based on the category name, please provide the most distinctive visual features unique to this category. Requirements: 1. Do not use the same appearance description as other categories. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid mentioning colors. Based on this category name, please provide the most unique feature that transcends visual concepts and distinguishes it from other categories, without visual descriptions. Requirements: 1. Do not use the same description as other categories. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid visual descriptions. Based on this category’s name, provide the most distinguishing feature description. Requirements: 1. Do not use the same description as other categories. 2. Keep it concise, no more than 10 words. 3. The description must only include visual and higher concept features of this category. 4. Ensure clarity, no ambiguity. 5. Do not describe color. 6. Must include visual feature of the category. 7. Must include higher concept feature of the category. a photo of a [CLASS].
Stl10 Based on the category name, please provide the most distinctive visual features unique to this category. Requirements: 1. Do not use the same appearance description as other categories. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid mentioning colors. Based on this category name, please provide the most unique feature that transcends visual concepts and distinguishes it from other categories, without visual descriptions. Requirements: 1. Do not use the same description as other categories. 2. Keep it concise, no more than 10 words. 3. Ensure clarity without ambiguity. 4. Avoid visual descriptions. Based on this category’s name, provide the most distinguishing feature description. Requirements: 1. Do not use the same description as other categories. 2. Keep it concise, no more than 10 words. 3. The description must only include visual and higher concept features of this category. 4. Ensure clarity, no ambiguity. 5. Do not describe color. 6. Must include visual feature of the category. 7. Must include higher concept feature of the category. a photo of a [CLASS].

### v3

@@ IV-C4 Position of Classname Varies
In the context of optimization, the position of the classname can be either at the end or in the middle [13]. Setting $\beta=0.6$, the context length is optimized to 16, and PK is used as the input. We explored the impact of different classname positions on the experimental results across 11 datasets. The outcomes are presented in Table IX. The overall performance is better when the classname position is at the end rather than in the middle.
IV-C5 Different Context Lengths
@@
We set the context lengths to 8, 16, and 32, with PK as the input and the $\beta$ value set to 0.6. We evaluated the mean comparison after training with 16 shots on 11 datasets to explore the impact of different context lengths on the results. The results are shown in Table X. Our method demonstrates robust performance across varying context lengths, with minimal perturbation to the outcomes attributable to changes in context length.
IV-C6 Longer Embeddings
To demonstrate that our method improves prediction performance through different types of multi-knowledge rather than by using longer embeddings, we employ three independent weight heads for contrastive learning. We then extend the length of the embeddings by concatenation for prediction. We use PK with ResNet50 as the backbone network, setting the $\beta$ value to 0.6. Three independent heads expand the embedding dimension from the original 1024 to 3072, represented by CoKnow-LE. The experimental results are presented in Table XI. The experimental results show that longer embeddings do not effectively improve prediction performance.
TABLE XI: Ablation Study of Longer Embeddings.
Dataset
Method Shots Caltech [35] DTD [37] EuroSAT [38] Aircraft [39] Flowers [41] Average
CoKnow-LE 1 $69.47$ $31.07$ $51.63$ $15.07$ $44.13$ $42.27$
2 $76.13$ $41.00$ $55.93$ $18.07$ $56.30$ $49.49$
4 $83.67$ $52.63$ $60.53$ $21.93$ $55.97$ $54.95$
8 $88.10$ $60.73$ $69.60$ $27.87$ $88.20$ $66.90$
16 $90.90$ $66.37$ $77.50$ $35.20$ $93.57$ $72.71$
CoKnow(ours) 1 89.30 47.93 58.67 18.97 75.83 58.14
2 90.72 51.45 66.97 22.02 82.63 62.76
4 90.90 58.23 74.60 25.67 88.67 67.61
8 91.40 62.90 79.83 30.47 92.77 71.47
16 92.83 67.20 84.77 37.57 95.33 93.10
V Limitations
This paper introduces Multi-Knowledge and effectively enhances the generalization performance of VLMs in downstream tasks. However, the ability of LLMs to generate the highest-quality Multi-Knowledge for maximizing the predictive power of VLMs has not yet been fully explored. As illustrated in Figure 5, using different LLMs requires further consideration of prompt strategies tailored to each model to fully leverage its potential. In the CoKnow framework proposed in this paper, the mapper layer employs a simple fully connected layer structure. As shown in Table IV, CoKnow does not achieve the highest performance metrics in the 1-shot setting, indicating that the generalization performance of this mapper structure needs improvement when the sample size is minimal. More advanced designs are needed in the future to address this issue.
VI Conclusion
@@
In this paper, we gain insights into the importance of multi-knowledge representation for VLMs. Existing technologies have overlooked the method of context template learning, which incorporates domain knowledge. Such oversight will restrict the ability of VLMs to learn multi-knowledge representations during the pretraining phase, further limiting its performance in downstream tasks. To address this issue, we propose context optimization with multi-knowledge representation (CoKnow). We trained lightweight semantic knowledge mappers capable of generating multi-knowledge representations without requiring additional inputs while also constructing a knowledge-guided prompt learning framework. We extensively evaluated it on multiple benchmark tests, and the results indicate that the proposed CoKnow is an effective method for prompt learning.